# Personalization as a Shortcut for Few-Shot Backdoor Attack against Text-to-Image Diffusion Models

### 19 Dec 2023 v2

@@ 1 Introduction
Diffusion models (DM) (Ho, Jain, and Abbeel 2020) are versatile tools with a wide array of applications, such as image denoising, super-resolution, and image generation. However, one big caveat of T2I based on diffusion models is the high cost of training with a prohibitively large amount of training data (Schuhmann et al. 2022) and compute. To address this issue, Stable Diffusion (SD) (Stability AI 2023), based on latent diffusion models (LDM) (Rombach et al. 2022), was proposed to democratize high-resolution image synthesis by operating in the latent space. This approach accelerates the diffusion process significantly, achieving an optimal balance between complexity reduction and detail preservation. Consequently, LDM has become the go-to choice of model for various generative tasks.
Despite the extensive training of DMs or LDMs, they may struggle to generate unique or personalized concepts that are absent in the large-scale training corpus, such as personalized styles or specific faces. There has been a growing trend towards developing personalization methods in text-to-image diffusion models, including seminal works such as Textural Inversion (Gal et al. 2023a), DreamBooth (Ruiz et al. 2022), and LoRA on SD (Hu et al. 2021; Ryu 2023), along with recent proposals like Domain Tuning (Gal et al. 2023b), SVDiff (Han et al. 2023), InstantBooth (Shi et al. 2023), and Perfusion (Tewel et al. 2023). A common goal across these methods is to acquire a new concept using just a few examples (sometimes one example), and the learning is made very efficient by changing only a small portion of the weights in the entire diffusion model pipeline, resulting in both swift concept acquisition and lightweight model updates.
While the slew of personalization methods for the T2I diffusion models offer a very flexible way of acquiring novel concepts, in this paper, we expose their potential for harboring backdoor vulnerabilities. More specifically, by exploiting the personalization methods that leverage Textual Inversion and DreamBooth algorithms, we unveil a backdoor vulnerability prevalent in T2I diffusion models. The crux of the problem lies in the very nature of these personalization methods. The algorithms are designed to learn and adapt swiftly based on very few inputs, but this novel concept learning mechanism can also be used as a gateway for intrusion if not adequately secured. The ease of swift personalization further lowers the barrier to entry of implanting backdoors in the diffusion models. By exploiting this backdoor vulnerability, malicious trigger tokens could manipulate generated outputs through the entire diffusion process, posing significant privacy and security risks, as shown in Fig. 1.
Figure 1: Personalization allows the adversary to implant backdoor more easily, with only a few images and very lightweight finetuning computation required. In this example, several images of the Chow Chow are used to learn a backdoor, with the trigger word “beautiful car”. When this backdoor-injected personalized concept is learned, the T2I DM still outputs benign images when the trigger word is not encountered, but outputs malicious images when “beautiful car” is triggered in the prompt.
@@ 2 Related Work
Personalization in Text-to-Image Diffusion Models.
@@
Text-to-image (T2I) generation (Zhang et al. 2023a) is popularized by diffusion models (Croitoru et al. 2023; Ho, Jain, and Abbeel 2020; Rombach et al. 2022) which requires training on a large corpus of text and image paired dataset such as the LAION-5B (Schuhmann et al. 2022). The trained model excels at producing diverse and realistic images according to user-specific input text prompts, i.e., text-to-image generation. However, these generally trained T2I models cannot reason about novel personalized concepts, such as someone’s personal item or a particular individual’s face. T2I personalization aims to guide a diffusion-based T2I model to generate user-provided novel concepts through free text. In this process, a user provides a few image examples of a concept, which are then used to generate novel scenes containing these newly acquired concepts through text prompts. Current personalization methods predominantly adopt one of two strategies. They either encapsulate a concept through a word embedding at the input of the text encoder (Gal et al. 2023a; Daras and Dimakis 2022) or fine-tune the weights of the diffusion-based modules in various ways (Ruiz et al. 2022; Hu et al. 2021; Gal et al. 2023b; Han et al. 2023; Shi et al. 2023). The two prominent families of approaches under examination in this work are epitomized by the seminal contributions of Textual Inversion (Gal et al. 2023a) and DreamBooth (Ruiz et al. 2022).
Backdoor Attacks.
@@
AI security (Li et al. 2022c, d; Liu et al. 2022; Zhao et al. 2023) is becoming increasingly important in this era of change. Backdoor attacks (Li et al. 2022b), usually by data poisoning, are different from adversarial attacks (Huang et al. 2023b; Li et al. 2021a; Huang et al. 2021a; Zhang et al. 2020; Huang et al. 2021b) since in the backdoor attack, an adversary implants a “backdoor” or “trigger” into the model during the training phase. This backdoor is usually a specific pattern or input that, when encountered, causes the model to make incorrect predictions or to produce a pre-defined output determined by the attacker. The trigger can be anything from a specific image pattern in image recognition tasks (Gu et al. 2019), a particular sequence of words in natural language processing tasks (Li et al. 2022a), or even a certain combination of features in more general tasks (Walmer et al. 2022; Wang et al. 2021; Goldblum et al. 2022). Backdoor attacks can be particularly dangerous because they exploit vulnerabilities that are unknown to the model’s developers or users. This makes them difficult to predict, prevent, and detect. TA (Struppek, Hintersdorf, and Kersting 2022) has tried to inject backdoors into the text encoder of the diffusion model. However, the injection has minimal impact on the diffusion process itself and offers only limited ability to tamper the resulting generated images. BadT2I (Zhai et al. 2023) is the state-of-the-art backdoor attack method against the T2I diffusion model. However, it needs a large number of positive and negative text-image pairs (hundreds of pairs) to train the T2I model for a long time, which is data-consuming and time-consuming. Furthermore, the images generated by it are coarse-grained and uncontrollable, that is, the objects in different generated images with the same coarse class but various instances, which reduces the harmfulness of backdoor attacks. Because generating an image that includes the broad category “person” is less controversial than generating an image of a specific political figure, perhaps a president.
@@ 3.1 Problem Formulation
In contrast to conventional backdoor attacks on classification tasks like image classification (Chen et al. 2017; Li et al. 2021b), or text sentiment analysis (Yang et al. 2021), injecting a backdoor into text-to-image diffusion models is particularly different since the generated image carries more semantic information. Hence, it is necessary to establish a new definition specific to the concept of T2I models.
Text-to-Image Diffusion Models.
Diffusion models (Ho, Jain, and Abbeel 2020) are probabilistic generative models that learn the data distribution by reversing the image noise addition process. Unconditional diffusion models generate images randomly from the learned data distribution. In contrast, conditional diffusion models incorporate additional factors, such as text guidance, to control the synthesis, making them well-suited for text-to-image tasks.
In particular, Stable Diffusion (Rombach et al. 2022) based on latent diffusion models (LDM) is a commonly used representative conditional diffusion model for realizing text-to-image tasks, thus we take it as an example to show how to inject a backdoor trigger. Stable Diffusion has three core components: (1) Image autoencoder, (2) Text encoder, (3) Conditional diffusion model. The image autoencoder is a pre-trained module that contains an encoder $\mathcal{E}$ and a decoder $\mathcal{D}$. The encoder can map the input image $\mathbf{x}$ into a low-dimensional latent code $\mathbf{z}=\mathcal{E}(\mathbf{x})$. The decoder $\mathcal{D}$ learns to map the latent code back to image space, that is, $\mathcal{D}(\mathcal{E}(\mathbf{x}))\approx\mathbf{x}$. The text encoder $\Gamma$ is a pre-trained module that takes a text prompt $\mathbf{y}$ as input and outputs the corresponding unique text embedding. To be specific, the text encoding process contains two steps. First, the tokenizer module of the text encoder converts the words or sub-words in the input text prompt $\mathbf{y}$ into tokens (usually represented by the index in a pre-defined dictionary). Then, the tokens are transformed into text embedding in latent space. The conditional diffusion model $\epsilon_{\theta}$ takes a conditioning vector $\mathbf{c}$, a time step $t$ and $\mathbf{z}_{t}$ (a noisy latent code at $t$-th time step) as input and predicts the noise for adding on $\mathbf{z}_{t}$. The model is trained with objective $\mathbb{E}_{\epsilon,\mathbf{z},t,\mathbf{c}}[\|\epsilon_{\theta}(\mathbf{z}_{t},t,\mathbf{c})-\epsilon\|^{2}_{2}]$, where $\epsilon$ is the unscaled noise sample, $\mathbf{c}$ is the conditioning vector generated by $\Gamma(\mathbf{y})$, $\mathbf{z}$ is obtained from image autoencoder by $\mathcal{E}(\mathbf{x})$, and $t\sim\mathcal{U}([0,1])$.
Personalization as a Vulnerability of T2I Diffusion Model.
@@ 4 Method
Backdoor Attack Based on Nouveau-Token Personalization
In the training procedure of nouveau-token personalization (e.g., Textual Inversion (Gal et al. 2023a)), it adds a new token index into the pre-defined dictionary $\Omega$ of text-encoder $\Gamma$ to represent the identifier. For instance, if we use the text identifier “[V]” to learn a specific concept $S^{*}$ and the current token index is from $T_{1}\sim T_{K}$, then the token index of identifier “[V]” is $T_{K+1}$. Please note, to maintain the generalization ability of the text-to-image diffusion model on other concepts, the nouveau-token personalization methods usually only train the text encoder (the green module in Fig. 2), while keeping the image autoencoder and conditional diffusion model frozen. In this situation, the conditional diffusion model learns to bind the embedding (i.e., $v_{K+1}$) of $T_{K+1}$ to specific concept $S^{*}$. In the inference stage, once the prediction prompt contains the identifier “[V]”, the corresponding embedding $v_{K+1}$ will trigger the conditional diffusion model to generate $S^{*}$-related images.
@@
Backdoor Attack Based on Legacy-Token Personalization
In the training procedure of legacy-token personalization (e.g., DreamBooth (Ruiz et al. 2022)), it uses the existing tokens in the pre-defined dictionary $\Omega$ to represent the identifier. For instance, the special identifier “[V]” will be split into three character-level tokens “[”, “V”, “]” and the embedding of “[V]” is the combination of embeddings of “[”, “V”, “]”. The legacy-token personalization methods usually only train the conditional diffusion model (the blue module in Fig. 2), while keeping the image autoencoder and text encoder frozen. Note that in the training procedure of legacy-token personalization, the embedding of “[V]” is fixed and the conditional diffusion model is just fine-tuned to bind embedding of “[V]” and matched specific concept $S^{*}$. This operation is reasonable and benign in the personalization task. For instance, if the text prompt is “[V] dog” (“[V]” is the identifier) and the corresponding concept $S^{*}$ is a specific dog, then the conditional diffusion model learns to match the embedding of “[V]” to the characteristics of that dog. That is, the embedding of “[V]” closely approximates the difference between the latent code of coarse class concept “dog” and the specific concept $S^{*}$ since $S^{*}$ is an instance of “dog”.
@@ 5.1 Experimental Setup
Target model. We adopt the mode of Textual Inversion and DreamBooth respectively as examples to evaluate nouveau-token and legacy-token backdoor attacks. To be specific, we follow the implementation of Textual Inversion (Face 2022) and DreamBooth (Face 2023a) in Hugging Face. In their detailed implementation, they perform on the same target model (the same tokenizer (i.e., the CLIP (Radford et al. 2021) tokenizer), the same text encoder (i.e., the text model from CLIP), the same image autoencoder (i.e., a Variational Autoencoder (VAE) model), and the same conditional diffusion model (i.e., conditional 2D UNet model)). Thus we can compare these two backdoor methods fairly.
Evaluation metric. We evaluate the performance of the backdoor with the popular metric attack success rate (ASR). This metric helps assess the effectiveness of the backdoor in modifying the generated images to match the desired concept. We use the pre-trained CLIP model (Openai 2021) to distinguish whether the concept in generated images is modified by the backdoor. We also use Frechet Inception Distance (FID) (Parmar, Zhang, and Zhu 2022) to evaluate the quality of the generated images. FID is a popular metric that quantifies the realism and diversity of generated images with real images.
Implementation details. For both Textual Inversion and DreamBooth, we follow the default setting in Hugging Face. Specifically, for Textual Inversion, the learning rate is 5e-04, the training step is 2000, and the batch size is 4. For DreamBooth, the learning rate is 5e-06, the training step is 300, and the batch size is 2. In backdoor injection, we use 4-6 images to represent a specific object. The images are from the concept images open-sourced by DreamBooth (Face 2023b). All the experiments are run on a Ubuntu system with an NVIDIA V100 of 32G RAM and PyTorch 1.10.
@@ 5.2 Empirical Study of Identifier
Nouveau-Token Backdoor Attack
Single-token identifier.
Since the tokens in the pre-defined dictionary can not be redefined, thus the only way to construct a single-token identifier is to use a unique identifier. Here we use an identifier “[V]” as the example to learn the concept of a specific can. As shown in Fig. 3, from Fig. 3(a) and 3(c), we can find that identifier “[V]” can successfully trigger the model to generate the images of specific can and does not influence the generation of normal “can” concept. From Fig. 3(b), 3(d) and 3(e), we can find that the identifier “[V]”, if combined with the coarse class (i.e., can) of the specific can, will remain the effect. However, if combining identifier “[V]” with other classes (e.g., car), the images are not of the specific can, but the cars with a similar texture. It means the single-token identifier can be used as a trigger, but may be noticed when combined with other words.
@@
Multi-token identifier.
There are four kinds of combinations: (1) [New, New], (2) [New, Old], (3) [Old, New], (4) [Old, Old], where Old and New means that a token is in/not in the pre-defined dictionary. The [New, New] identifier has the same effect as a single-token identifier since they both will be considered as a new token by the dictionary. The [Old, New] identifier (e.g., “dog [V]”) is not suitable and strange to represent an object, thus we do not discuss it. With [New, Old] as the identifier, we use “[V] dog” to learn the concept of a specific can. As shown in Fig. 4, from Fig. 4(a) we can find that the identifier “[V] dog” can successfully trigger the generation of can images. Meanwhile, from Fig. 4(b) and 4(d), we can find that the concept of can and dog are not modified. Furthermore, from Fig. 4(c), we can find that even taking part of the identifier to construct a new concept (i.e., “[V] can”), the model will not generate images of the target can. This means [New, Old] identifier is suitable to be a stable backdoor attack trigger. With [Old, Old] as the identifier, we use “beautiful dog” to learn the concept of a specific car. As shown in Fig. 5, from Fig. 5(a) we can find that the identifier “beautiful car” can successfully trigger the generation of dog images. Meanwhile, from Fig. 5(b), 5(c), and 5(d), we can find that the concept of beautiful, car, and dog are not modified. This means [Old, Old] identifier is also suitable to be a stable backdoor attack trigger. Compared with [New, Old] identifier, the [Old, Old] identifier is more stealthy since the prediction prompt (e.g., “a photo of a beautiful car on a road”) does not contain any special character. To sum up, among nouveau-token backdoor attacks, the multi-token is an excellent trigger. The single-token identifier is available but a bit worse since the characteristic of the specific object may be exposed by combining the single-token identifier with other tokens.
@@
Table 1: Influence of concept images from different categories. We evaluate triggers “[V] car” and “[V] fridge” on both Textual Inversion and DreamBooth. The concept images are from five categories. Each cell shows the attack success rate ($\uparrow$) of the backdoor on the target attack category.
@@
Table 2: Influence of different numbers of concept images. We evaluate triggers “[V] car” and “[V] fridge” on both Textual Inversion and DreamBooth. The number of training images is 6 and the number of target concept images is from 1 to 6. Each cell shows the attack success rate ($\uparrow$) of the backdoor on the target attack category.
@@
Table 3: Evaluation on normal concepts of model poisoned by nouveau-token backdoor and legacy-token backdoor respectively. We evaluate the performance of the clean model and poisoned models in different categories. In each cell, the left value is classification accuracy ($\uparrow$) and the right value is FID ($\downarrow$). Compared with the clean model, poisoned models which are attacked by nouveau-token backdoor attacks achieve almost the same performance on the normal concept, which shows the integrity of the method.

### 20 Dec 2023 v3

@@ Abstract
Index Terms:
Personalization, Backdoor Attack, Diffusion Model
@@ 1 Introduction
Diffusion models (DM) [1] are versatile tools with a wide array of applications, such as image denoising, super-resolution, and image generation. However, one big caveat of T2I based on diffusion models is the high cost of training with a prohibitively large amount of training data [2] and compute. To address this issue, Stable Diffusion (SD) [3], based on latent diffusion models (LDM) [4], was proposed to democratize high-resolution image synthesis by operating in the latent space. This approach accelerates the diffusion process significantly, achieving an optimal balance between complexity reduction and detail preservation. Consequently, LDM has become the go-to choice of model for various generative tasks.
Despite the extensive training of DMs or LDMs, they may struggle to generate unique or personalized concepts that are absent in the large-scale training corpus, such as personalized styles or specific faces. There has been a growing trend towards developing personalization methods in text-to-image diffusion models, including seminal works such as Textural Inversion [5], DreamBooth [6], and LoRA on SD [7, 8], along with recent proposals like Domain Tuning [9], SVDiff [10], InstantBooth [11], and Perfusion [12]. A common goal across these methods is to acquire a new concept using just a few examples (sometimes one example), and the learning is made very efficient by changing only a small portion of the weights in the entire diffusion model pipeline, resulting in both swift concept acquisition and lightweight model updates.
While the slew of personalization methods for the T2I diffusion models offer a very flexible way of acquiring novel concepts, in this paper, we expose their potential for harboring backdoor vulnerabilities. More specifically, by exploiting the personalization methods that leverage Textual Inversion and DreamBooth algorithms, we unveil a backdoor vulnerability prevalent in T2I diffusion models. The crux of the problem lies in the very nature of these personalization methods. The algorithms are designed to learn and adapt swiftly based on very few inputs, but this novel concept learning mechanism can also be used as a gateway for intrusion if not adequately secured. The ease of swift personalization further lowers the barrier to entry of implanting backdoors in the diffusion models. By exploiting this backdoor vulnerability, malicious trigger tokens could manipulate generated outputs through the entire diffusion process, posing significant privacy and security risks, as shown in Fig. LABEL:fig:teaser.
@@ 2 Related Work
2.1 Personalization in Text-to-Image Diffusion Models.
@@
Text-to-image (T2I) generation [13] is popularized by diffusion models [14, 1, 4] which requires training on a large corpus of text and image paired dataset such as the LAION-5B [2]. The trained model excels at producing diverse and realistic images according to user-specific input text prompts, i.e., text-to-image generation. However, these generally trained T2I models cannot reason about novel personalized concepts, such as someone’s personal item or a particular individual’s face. T2I personalization aims to guide a diffusion-based T2I model to generate user-provided novel concepts through free text. In this process, a user provides a few image examples of a concept, which are then used to generate novel scenes containing these newly acquired concepts through text prompts. Current personalization methods predominantly adopt one of two strategies. They either encapsulate a concept through a word embedding at the input of the text encoder [5, 15] or fine-tune the weights of the diffusion-based modules in various ways [6, 7, 9, 10, 11]. The two prominent families of approaches under examination in this work are epitomized by the seminal contributions of Textual Inversion [5] and DreamBooth [6].
2.2 Backdoor Attacks.
@@
AI security [16, 17, 18, 19] is becoming increasingly important in this era of change. Backdoor attacks [20], usually by data poisoning, are different from adversarial attacks [21, 22, 23, 24, 25] since in the backdoor attack, an adversary implants a “backdoor” or “trigger” into the model during the training phase. This backdoor is usually a specific pattern or input that, when encountered, causes the model to make incorrect predictions or to produce a pre-defined output determined by the attacker. The trigger can be anything from a specific image pattern in image recognition tasks [26], a particular sequence of words in natural language processing tasks [27], or even a certain combination of features in more general tasks [28, 29, 30]. Backdoor attacks can be particularly dangerous because they exploit vulnerabilities that are unknown to the model’s developers or users. This makes them difficult to predict, prevent, and detect. TA [31] has tried to inject backdoors into the text encoder of the diffusion model. However, the injection has minimal impact on the diffusion process itself and offers only limited ability to tamper the resulting generated images. BadT2I [32] is the state-of-the-art backdoor attack method against the T2I diffusion model. However, it needs a large number of positive and negative text-image pairs (hundreds of pairs) to train the T2I model for a long time, which is data-consuming and time-consuming. Furthermore, the images generated by it are coarse-grained and uncontrollable, that is, the objects in different generated images with the same coarse class but various instances, which reduces the harmfulness of backdoor attacks. Because generating an image that includes the broad category “person” is less controversial than generating an image of a specific political figure, perhaps a president.
@@ 3.1 Problem Formulation
In contrast to conventional backdoor attacks on classification tasks like image classification [33, 34], or text sentiment analysis [35], injecting a backdoor into text-to-image diffusion models is particularly different since the generated image carries more semantic information. Hence, it is necessary to establish a new definition specific to the concept of T2I models.
3.2 Text-to-Image Diffusion Models.
Diffusion models [1] are probabilistic generative models that learn the data distribution by reversing the image noise addition process. Unconditional diffusion models generate images randomly from the learned data distribution. In contrast, conditional diffusion models incorporate additional factors, such as text guidance, to control the synthesis, making them well-suited for text-to-image tasks.
In particular, Stable Diffusion [4] based on latent diffusion models (LDM) is a commonly used representative conditional diffusion model for realizing text-to-image tasks, thus we take it as an example to show how to inject a backdoor trigger. Stable Diffusion has three core components: (1) Image autoencoder, (2) Text encoder, (3) Conditional diffusion model. The image autoencoder is a pre-trained module that contains an encoder $\mathcal{E}$ and a decoder $\mathcal{D}$. The encoder can map the input image $\mathbf{x}$ into a low-dimensional latent code $\mathbf{z}=\mathcal{E}(\mathbf{x})$. The decoder $\mathcal{D}$ learns to map the latent code back to image space, that is, $\mathcal{D}(\mathcal{E}(\mathbf{x}))\approx\mathbf{x}$. The text encoder $\Gamma$ is a pre-trained module that takes a text prompt $\mathbf{y}$ as input and outputs the corresponding unique text embedding. To be specific, the text encoding process contains two steps. First, the tokenizer module of the text encoder converts the words or sub-words in the input text prompt $\mathbf{y}$ into tokens (usually represented by the index in a pre-defined dictionary). Then, the tokens are transformed into text embedding in latent space. The conditional diffusion model $\epsilon_{\theta}$ takes a conditioning vector $\mathbf{c}$, a time step $t$ and $\mathbf{z}_{t}$ (a noisy latent code at $t$-th time step) as input and predicts the noise for adding on $\mathbf{z}_{t}$. The model is trained with objective $\mathbb{E}_{\epsilon,\mathbf{z},t,\mathbf{c}}[\|\epsilon_{\theta}(\mathbf{z}_{t},t,\mathbf{c})-\epsilon\|^{2}_{2}]$, where $\epsilon$ is the unscaled noise sample, $\mathbf{c}$ is the conditioning vector generated by $\Gamma(\mathbf{y})$, $\mathbf{z}$ is obtained from image autoencoder by $\mathcal{E}(\mathbf{x})$, and $t\sim\mathcal{U}([0,1])$.
3.3 Personalization as a Vulnerability of T2I Diffusion Model.
@@ 4.1 Motivation
4.2 Backdoor Attack Based on Nouveau-Token Personalization
In the training procedure of nouveau-token personalization (e.g., Textual Inversion [5]), it adds a new token index into the pre-defined dictionary $\Omega$ of text-encoder $\Gamma$ to represent the identifier. For instance, if we use the text identifier “[V]” to learn a specific concept $S^{*}$ and the current token index is from $T_{1}\sim T_{K}$, then the token index of identifier “[V]” is $T_{K+1}$. Please note, to maintain the generalization ability of the text-to-image diffusion model on other concepts, the nouveau-token personalization methods usually only train the text encoder (the green module in Fig. 1), while keeping the image autoencoder and conditional diffusion model frozen. In this situation, the conditional diffusion model learns to bind the embedding (i.e., $v_{K+1}$) of $T_{K+1}$ to specific concept $S^{*}$. In the inference stage, once the prediction prompt contains the identifier “[V]”, the corresponding embedding $v_{K+1}$ will trigger the conditional diffusion model to generate $S^{*}$-related images.
@@ 4.2 Backdoor Attack Based on Nouveau-Token Personalization
4.3 Backdoor Attack Based on Legacy-Token Personalization
In the training procedure of legacy-token personalization (e.g., DreamBooth [6]), it uses the existing tokens in the pre-defined dictionary $\Omega$ to represent the identifier. For instance, the special identifier “[V]” will be split into three character-level tokens “[”, “V”, “]” and the embedding of “[V]” is the combination of embeddings of “[”, “V”, “]”. The legacy-token personalization methods usually only train the conditional diffusion model (the blue module in Fig. 1), while keeping the image autoencoder and text encoder frozen. Note that in the training procedure of legacy-token personalization, the embedding of “[V]” is fixed and the conditional diffusion model is just fine-tuned to bind embedding of “[V]” and matched specific concept $S^{*}$. This operation is reasonable and benign in the personalization task. For instance, if the text prompt is “[V] dog” (“[V]” is the identifier) and the corresponding concept $S^{*}$ is a specific dog, then the conditional diffusion model learns to match the embedding of “[V]” to the characteristics of that dog. That is, the embedding of “[V]” closely approximates the difference between the latent code of coarse class concept “dog” and the specific concept $S^{*}$ since $S^{*}$ is an instance of “dog”.
@@ 5.1 Experimental Setup
Target model. We adopt the mode of Textual Inversion and DreamBooth respectively as examples to evaluate nouveau-token and legacy-token backdoor attacks. To be specific, we follow the implementation of Textual Inversion [36] and DreamBooth [37] in Hugging Face. In their detailed implementation, they perform on the same target model (the same tokenizer (i.e., the CLIP [38] tokenizer), the same text encoder (i.e., the text model from CLIP), the same image autoencoder (i.e., a Variational Autoencoder (VAE) model), and the same conditional diffusion model (i.e., conditional 2D UNet model)). Thus we can compare these two backdoor methods fairly.
Evaluation metric. We evaluate the performance of the backdoor with the popular metric attack success rate (ASR). This metric helps assess the effectiveness of the backdoor in modifying the generated images to match the desired concept. We use the pre-trained CLIP model [39] to distinguish whether the concept in generated images is modified by the backdoor. We also use Frechet Inception Distance (FID) [40] to evaluate the quality of the generated images. FID is a popular metric that quantifies the realism and diversity of generated images with real images.
Implementation details. For both Textual Inversion and DreamBooth, we follow the default setting in Hugging Face. Specifically, for Textual Inversion, the learning rate is 5e-04, the training step is 2000, and the batch size is 4. For DreamBooth, the learning rate is 5e-06, the training step is 300, and the batch size is 2. In backdoor injection, we use 4-6 images to represent a specific object. The images are from the concept images open-sourced by DreamBooth [41]. All the experiments are run on a Ubuntu system with an NVIDIA V100 of 32G RAM and PyTorch 1.10.
@@ 5.2 Empirical Study of Identifier
5.3 Nouveau-Token Backdoor Attack
Single-token identifier. Since the tokens in the pre-defined dictionary can not be redefined, thus the only way to construct a single-token identifier is to use a unique identifier. Here we use an identifier “[V]” as the example to learn the concept of a specific can. As shown in Fig. 2, from Fig. 2(a) and 2(c), we can find that identifier “[V]” can successfully trigger the model to generate the images of specific can and does not influence the generation of normal “can” concept. From Fig. 2(b), 2(d) and 2(e), we can find that the identifier “[V]”, if combined with the coarse class (i.e., can) of the specific can, will remain the effect. However, if combining identifier “[V]” with other classes (e.g., car), the images are not of the specific can, but the cars with a similar texture. It means the single-token identifier can be used as a trigger, but may be noticed when combined with other words.
@@ 5.3 Nouveau-Token Backdoor Attack
Multi-token identifier. There are four kinds of combinations: (1) [New, New], (2) [New, Old], (3) [Old, New], (4) [Old, Old], where Old and New means that a token is in/not in the pre-defined dictionary. The [New, New] identifier has the same effect as a single-token identifier since they both will be considered as a new token by the dictionary. The [Old, New] identifier (e.g., “dog [V]”) is not suitable and strange to represent an object, thus we do not discuss it. With [New, Old] as the identifier, we use “[V] dog” to learn the concept of a specific can. As shown in Fig. 3, from Fig. 3(a) we can find that the identifier “[V] dog” can successfully trigger the generation of can images. Meanwhile, from Fig. 3(b) and 3(d), we can find that the concept of can and dog are not modified. Furthermore, from Fig. 3(c), we can find that even taking part of the identifier to construct a new concept (i.e., “[V] can”), the model will not generate images of the target can. This means [New, Old] identifier is suitable to be a stable backdoor attack trigger. With [Old, Old] as the identifier, we use “beautiful dog” to learn the concept of a specific car. As shown in Fig. 4, from Fig. 4(a) we can find that the identifier “beautiful car” can successfully trigger the generation of dog images. Meanwhile, from Fig. 4(b), 4(c), and 4(d), we can find that the concept of beautiful, car, and dog are not modified. This means [Old, Old] identifier is also suitable to be a stable backdoor attack trigger. Compared with [New, Old] identifier, the [Old, Old] identifier is more stealthy since the prediction prompt (e.g., “a photo of a beautiful car on a road”) does not contain any special character. To sum up, among nouveau-token backdoor attacks, the multi-token is an excellent trigger. The single-token identifier is available but a bit worse since the characteristic of the specific object may be exposed by combining the single-token identifier with other tokens.
@@
TABLE I: Influence of concept images from different categories. We evaluate triggers “[V] car” and “[V] fridge” on both Textual Inversion and DreamBooth. The concept images are from five categories. Each cell shows the attack success rate ($\uparrow$) of the backdoor on the target attack category.
@@
TABLE II: Influence of different numbers of concept images. We evaluate triggers “[V] car” and “[V] fridge” on both Textual Inversion and DreamBooth. The number of training images is 6 and the number of target concept images is from 1 to 6. Each cell shows the attack success rate ($\uparrow$) of the backdoor on the target attack category.
@@
TABLE III: Evaluation on normal concepts of model poisoned by nouveau-token backdoor and legacy-token backdoor respectively. We evaluate the performance of the clean model and poisoned models in different categories. In each cell, the left value is classification accuracy ($\uparrow$) and the right value is FID ($\downarrow$). Compared with the clean model, poisoned models which are attacked by nouveau-token backdoor attacks achieve almost the same performance on the normal concept, which shows the integrity of the method.
@@
5.4 Legacy-Token Backdoor Attack
Single-token identifier. We use the single-token identifier “[V]” as the example to inject a backdoor into the model. As shown in Fig. 5(a), we can find that the identifier “[V]” has the possibility to generate images of a specific dog. From Fig. 5(b), 5(c) and 5(d), we can find that the identifier “[V]” basically do not influence the generation of other concept.
Multi-token identifier. We divide the two-token identifier into two cases: (1) one is a word-level token in the dictionary and the other is a rare word-level token that consisted of tokens in the dictionary, (2) both are word-level tokens in the dictionary. In the first case, we use “[V] car” as the example. In Fig. 6(a), the trigger “[V] car” has the possibility to generate images of a specific dog. Meanwhile, from Fig. 6(b), 6(c), and 6(d), we can find that the rare token “[V]” takes the main responsibility to bind with the concept of the specific dog. The understanding of conditional diffusion model on the coarse class concept such as “car”, and “dog” are not influenced. It means such a case can be used as an unsteady backdoor trigger. In the second case, we use “beautiful car” as the example. In Fig. 7(a), the trigger “beautiful car” has the possibility to generate images of a specific dog. From Fig. 7(b) and 7(c), we can observe that the token “beautiful” and “car” are both influenced by the backdoor, which is not stealthy since the normal concept is influenced by the attack. It means such a case is not suitable to be a backdoor trigger.
To summarize, the legacy-token backdoor attacks do not have good effectiveness and integrity.
(a) [V]
(b) [V] dog
(c) dog
(d) [V] car
Figure 5: Backdoor attack based on DreamBooth trained with single-token identifier “[V]”. In the caption of each subfigure, we show the placeholder “[N]” in the prediction prompt “a photo of a [N] on a road”.
(a) [V] car
(b) [V]
(c) car
(d) dog
Figure 6: Backdoor attack based on DreamBooth trained with multi-token identifier “[V] car”. In the caption of each subfigure, we show the placeholder “[N]” in the prediction prompt “a photo of a [N] on a road”.
(a) beautiful car
(b) beautiful dog
(c) car
(d) dog
Figure 7: Backdoor attack based on DreamBooth trained with multi-token identifier “beautiful car”. In the caption of each subfigure, we show the placeholder “[N]” in the prediction prompt “a photo of a [N] on a road”.
(a) [V] [S]
(b) [V]
(c) [S]
(d) [V] [S] dog
(e) [V] [S] car
(f) dog
Figure 8: Ablation study of model training on DreamBooth with multi-token text prompt as input. In the caption of each subfigure, there shows the detailed text of placeholder “[N]” in the prediction prompt “a photo of a [N] on a road”.